A 10 kg block is moving at 6 m/s, it is slowed to a speed of 2 m/s in 15 s. How much force is required for this to occur?



1
Expert's answer
2021-01-26T08:43:27-0500

Find the acceleration:


"a=\frac{v_f-v_i}{t}=\frac{2-6}{15}=-0.267\text{ m/s}^2."

Thus, according to Newton's second law, the force is


"F=ma=-2.67\text{ N}."

The force is negative because it slows down the block.